# Adaptive Circuit Behavior and Generalization in Mechanistic Interpretability

### v1

@@ 3.2 Base IOI Circuit Performance on Variants
3.3 Circuit Functionality Analysis
The sharp deviation in performance between the base IOI circuit and the full model raises an important question: Are the elements of the circuit maintaining the same functions as they had on the base IOI task, or have their functions changed in response to changes in the prompt? To understand this further, we calculate the differences between the attention scores of all heads in the base IOI circuit on the base IOI and DoubleIO datasets. Since the full model may respond differently to the prompt variants, we do the same calculations for the model as well. The results are shown in Figure 2.
@@ 3.3 Circuit Functionality Analysis
Overall, these results suggest that most components of the base IOI circuit maintain their original functionality even for the prompt variants. We investigate the causes of this behavior in Section 4.
@@ 4 S2 Hacking: Performance without Faithfulness
As established in the previous section, the base IOI circuit significantly outperforms the full model on the IOI prompt variants without any changes in the behavior of its components, which appears to be unfaithful to the full model. In this section, we identify the source of the deviations in behavior and functionality between the base IOI circuit and the full model.
The basis of the discrepancy between the circuit and the model performance is a mechanism in the base IOI circuit which we term S2 Hacking. In the base IOI circuit, the Induction and Duplicate Token heads primarily attend to the S2 token, which is always the incorrect answer in each of the IOI prompt variants. The outputs of these heads are then used by the S-Inhibition heads to suppress the attention on the S1 and S2 tokens. This mechanism is a key factor in how the circuit is able to return the correct (IO) token with high probability.
@@
This knockout procedure effectively points the circuit toward the correct answer every time, since the S2 token is the only input to the Duplicate Token heads that is not knocked out. These heads feed into the Induction and S-Inhibition heads, which are directed to only attend to the S2 token. As a result, they always suppress attention on the subject tokens (S1, S2), which pushes the Name Mover heads towards returning an IO token. This mechanism of consistently suppressing the subject tokens and returning the IO token enables the base IOI circuit to solve the task across all of the prompt variants.
We refer to this phenomenon as S2 Hacking. Note that this phenomenon only occurs in the base IOI circuit, as it is a byproduct of the knockout procedure for evaluating the circuit and not actually how the full model solves the task. This is evident from how much the circuit performance deviates from the model performance. Additionally, S2 Hacking is not observed on base IOI inputs since they only have one duplicated name (S), and was discovered by generalizing the prompt format. For the rest of this section, we focus on the DoubleIO prompt variant and demonstrate how S2 Hacking occurs within the base IOI circuit.
@@ 4.2 Tracing the S2 Hacking Mechanism
Our results demonstrate that the S2 Hacking mechanism is primarily carried out through S-Inhibition head 8.6, Induction heads 5.9 and 5.5, and Duplicate head 3.0. Through this mechanism, we find that all of the Name Mover heads show higher confidence in predicting an IO token over an S token. The remaining heads in the circuit show similar behavior to their performance on base IOI prompts. Additional details on these experiments can be found in Appendix A.
@@ Appendix A S2 Hacking
A.1 S-Inhibition Heads
@@ A.3 Duplicate Heads
A.4 Previous Token Heads
@@ Appendix B Identifying Layer 2 Head 2 As The Decision Point
Figure 9: Attention scores of Layer 4 Head 11 indicate that the Previous Token Head attends equally to both the IO and S previous tokens
@@
(a) Attention (IO first): Layer 2, Head 2. The attention patterns favor IO when IO appears first.
(b) Attention (Subject first): Layer 2, Head 2. The attention patterns favor Subject when Subject appears first.
@@
Figure 11: Here we show the difference between the absolute direct effects between the S2 and IO2 edges of the same node. The high value of nodes provide preliminary evidence that they might be a decision point.
@@
Figure 12: Change in direct effect from subject to IO token for induction head and duplicate token heads .

### v2

@@ 3.2 Base IOI Circuit Performance on Variants
3.3 Head-Level Attention Patterns on Base IOI and Variants
The sharp deviation in performance between the base IOI circuit and the full model raises an important question: Are the elements of the circuit maintaining the same functions as they had on the base IOI task, or have their functions changed in response to changes in the prompt? To examine this further, we calculate the average difference between the attention scores of each head in the base IOI circuit for its most relevant token position (specified in Wang et al. (2023)) on the base IOI and prompt variant datasets. Since the full model may respond differently to the prompt variants, we do the same calculations for these heads in the model as well. The results are shown in Figure 2.
@@ 3.3 Head-Level Attention Patterns on Base IOI and Variants
Overall, these results suggest that most components of the base IOI circuit show lower deviations in attention patterns on base IOI inputs and the prompt variants compared to the full model. We later show in Section 5 that these heads retain their original functionalities from the base IOI circuit.
@@ 4 S2 Hacking: Performance without Faithfulness
We observed in the previous section that the base IOI circuit significantly outperforms the full model on the IOI prompt variants, with most of its components maintaining similar attention patterns to base IOI inputs. In this section, we identify the source of the deviations in behavior and performance between the base IOI circuit and the full model.
The basis of the discrepancy between the circuit and the model performance is a mechanism in the base IOI circuit which we term S2 Hacking. In the base IOI circuit, the Induction and Duplicate Token heads are primarily active at the S2 token, which is always the incorrect answer in each of the IOI prompt variants. The outputs of these heads are then used by the S-Inhibition heads to suppress the attention on the S1 and S2 tokens. This mechanism is a key factor in how the circuit is able to return the correct (IO) token with high probability.
@@
This knockout procedure effectively points the circuit toward the correct answer every time, since the S2 token is the only input to the Duplicate Token heads that is not knocked out. These heads feed into the Induction and S-Inhibition heads, which are directed to only attend to subject-related tokens, S+1 and S2 respectively. As a result, the S-Inhibition heads always suppress attention on the subject tokens (S1, S2), which pushes the Name Mover heads towards returning an IO token. This mechanism of consistently suppressing the subject tokens and returning the IO token enables the base IOI circuit to solve the task across all of the prompt variants.
We refer to this phenomenon as S2 Hacking. Note that this phenomenon only occurs in the base IOI circuit, as it is a byproduct of the knockout procedure for evaluating the circuit and not actually how the full model solves the task. This is evident from how much the circuit performance deviates from the model performance. Additionally, S2 Hacking is not observed on base IOI inputs since they only have one duplicated name (S), and was discovered by generalizing the prompt format. For the rest of this section, we focus on the DoubleIO prompt variant and demonstrate how S2 Hacking occurs within the base IOI circuit. Additional details on all metrics and experiments are in Appendix A.
@@ 4.2 Tracing the S2 Hacking Mechanism
Our results demonstrate that the S2 Hacking mechanism is primarily carried out through S-Inhibition head 8.6, Induction heads 5.9 and 5.5, and Duplicate head 3.0. Through this mechanism, we find that all of the Name Mover heads show higher confidence in predicting an IO token over an S token. The remaining heads in the circuit show similar behavior to their performance on base IOI prompts.
@@ Appendix A S2 Hacking
A.1 Metrics
Each head type is defined by a query position (“active” token) and a key position (“attended” token). For DoubleIO prompts, most heads have a decision to focus on tokens leading to the prediction of either S or IO. For each head type, Wang et al. (2023) specify which tokens to primarily attend to in order to produce the correct answer. We refer to these as the “correct” and “incorrect” tokens for that head type, and they are used in the definition of the confidence ratio in Section 4.1. We specify these “correct” and “incorrect” tokens for each head type below.
For Inhibition heads, which are active at END, attending to S2 suppresses the prediction of the subject token and increases the probability of the Name Mover heads returning the IO token, which is the “correct” output. Attending to IO2, which suppresses the prediction of the IO token, is considered “incorrect”. We refer to the attention probability of S2 at END position as the “correct” probability, while the incorrect probability is the attention probability of IO2 at END. Similarly, for Induction heads we refer to the attention probability of S1+1 at S2 as “correct”, and the attention probability of IO1+1 at IO2 as “incorrect”. For Duplicate Token heads, we refer to the attention probability of S1 at S1+1 as “correct”, and the attention probability of IO1 at IO1+1 as “incorrect”. For Previous Token heads, we refer to the attention probability of S1 at S1+1 as “correct”, and the attention probability of IO1 at IO1+1 as “incorrect”.
A.2 S-Inhibition Heads
@@ A.4 Duplicate Heads
A.5 Previous Token Heads
@@ Appendix B Identifying Layer 2 Head 2 As The Decision Point
Figure 9: Attention scores of Layer 4 Head 11 indicate that the Previous Token Head attends equally to both the IO and S previous tokens.
@@
(a) Attention (IO first): The attention patterns favor IO when IO appears first.
(b) Attention (Subject first): The attention patterns favor Subject when Subject appears first.
@@
Figure 11: Difference between the absolute direct effects between the S2 and IO2 edges of the same node. The high value of nodes provide preliminary evidence that they might be a decision point.
@@
Figure 13: Change in direct effect from subject to IO token for Induction and Duplicate Token heads.